# Logics-STEM: Empowering LLM Reasoning via Failure-Driven Post-Training and Document Knowledge Enhancement

### v1

@@ 1 Introduction
In the past years, large language models (LLMs) like OpenAI’s o1 series (jaech2024openai), QwQ (qwq32b), and DeepSeek-R1 (guo2025deepseek) have demonstrated strong performance on challenging reasoning tasks in the field of mathematics and broader STEM. The reasoning capabilities in these models typically emerge from post-training techniques such as supervised fine-tuning (SFT) and/or reinforcement learning (RL) based on strong foundation models. However, while many models are open-sourced, the underlying post-training pipeline and training data curation remain undisclosed, creating both challenges and opportunities for future work.
More recently, the open-source community has made substantial efforts to develop data-construction recipes and algorithmic strategies for cultivating advanced reasoning capabilities in small-scale models, leading to a series of notable works such as Klear-Reasoner (su2025klear), Ring-Lite (DBLP:journals/corr/abs-2506-14731), MiMo (coreteam2025mimounlockingreasoningpotential), OpenThoughts (guha2025openthoughts), Llama-Nemotron (bercovich2025llama), and AceReason-Nemotron liu2025acereason, among others (fu2025areal; deepscaler2025; DBLP:journals/corr/abs-2503-10460). However, despite these empirical successes, the community still lacks a unified framework to guide data curation and exploitation through post-training algorithms. It is widely recognized in the LLM community that “data is the new oil”, and the algorithm can only succeed when it effectively captures the desired data distribution. This motivates a central question in training reasoning models:
@@
In this report, we address this question from both theoretical and engineering perspectives. We first provide a data-centric view of the widely adopted SFT-RL pipeline (ouyang2022training) by framing it as a distribution-matching problem. We hypothesize that the first-stage SFT builds a strong proposal distribution that one draws samples for following usage; and the second stage post-training, no matter SFT or RL, shifts the model towards a gold-standard target distribution behind desirable properties such as reasoning ability.
@@
Specifically, we curate reasoning data from publicly available datasets and further augment it with synthetic examples generated from documents parsed by Logics-Parsing (chen2025logicsparsingtechnicalreport). Coupled with a fine-grained, difficulty-aware stratified sampling strategy, extensive experiments show that our curated Logics-STEM-SFT-Dataset already equips LLMs with strong foundational reasoning capabilities. Moreover, we adopt a failure-driven post-training paradigm to further align the model with the gold-standard reasoning distribution. Concretely, after the first-stage SFT, we perform targeted knowledge retrieval and synthesize data around the model’s failure regions to guide a second-stage SFT or RL. This yields two alternative pipelines: SFT–RL and SFT–SFT. We systematically test our method under both SFT-RL and SFT-SFT pipelines and show that our approach substantially improves the model’s reasoning ability.
@@ 2 SFT-RL pipeline as a distribution matching
After pre-training, SFT followed by RL has become a widely adopted recipe for improving LLMs’ reasoning ability (ouyang2022training). SFT is typically used to familiarize the model with long chain-of-thought (CoT) reasoning traces, while RL further aligns the model with human preferences or sharpens the policy distribution to produce more satisfactory responses with fewer samples (yue2025does; wen2025reinforcementlearningverifiablerewards). From this perspective, the overall post-training procedure can be viewed as optimizing an expected objective estimated via Monte Carlo sampling. Let the training data consist of pairs $(x,y)$, where $x$ is the prompt and $y$ is the target output (e.g., a long-CoT response ending with a boxed final answer). Denote by $y^{t}$ the $t$-th token in yy [^1] We use subscripts $y_{i}$ to denote the $i$-th training example and superscripts $y^{t}$ to denote the $t$-th token within an example.. We write the sample-level supervised loss as $\ell_{\theta}(x,y)$ (e.g., the negative log-likelihood, NLL), where $\theta$ are the model parameters. The training objective is then,
@@
We elaborate on the remark a bit. SFT fits the model to $P_{0}$ through a large-scale dataset ${\mathcal{D}}\sim P_{0}$, which provides good proposal distribution with broad coverage and grants the model with general reasoning abilities. However, at this stage, the mismatch between $P_{0}$ and $P^{*}$ would make the optimization biased, thus performing suboptimally on specific tasks. The second stage training is used to implicitly eliminate the bias. For instance, a vanilla policy gradient can be considered as fitting a distribution depending on the advantage function $A(x,y)$ (DBLP:conf/icml/0011C00LZ23),
@@ 3.1 Data Curation Pipeline
Previous studies (guha2025openthoughts; liu2025acereason; openr1) have demonstrated the effectiveness of open-source data. To fully leverage open-source data and ensure diversity, we aggregate questions from a wide range of highly regarded and frequently cited sources on Huggingface [^2] https://huggingface.co/, as listed in Appendix A.1. For the purpose of preventing potential quality risks, all early synthetic subsets and multi-modal subsets are excluded. Additional, we utilize DEGISNER (liu2025designer) to incorporate synthetic datasets from books and webpages (DLR-Book and DLR-Web in liu2025designer) to further scale dataset volume and enhance diversity. All PDF-format documents are parsed into plain text by Logics-Parsing (chen2025logicsparsingtechnicalreport) and subsequently processed through the DESIGNER synthesis pipeline.
@@
To balance computational efficiency and response accuracy, Qwen3-235B-A22B-Thinking-2507444https://huggingface.co/Qwen/Qwen3-235B-A22B-Thinking-2507 serves as the teacher model to distill reasoning responses for each question. Generation configurations are detailed in Appendix A.3. To suppress systematic bias during subsequent training, we discard responses that exhibit degenerative repetition. Concretely, any response with an n-gram duplication ratio exceeding a predefined threshold is excluded from the training dataset. Optional further answer verification is conducted using math-verify library [^5] https://github.com/huggingface/Math-Verify. In cases where responses do not match the standard verifiable answers, the teacher model is employed to regenerate the response once more. The strategy of filtering out incorrect responses is not adopted, as empirical results indicate that it leads to degraded performance, consistent with the conclusion of OpenThoughts (guha2025openthoughts).
@@ 3.2 Weighted Stratified Sampling
Despite the implementation of extensive filtering strategies, the volume of data generated by our pipeline remains a significant challenge for model training, thereby necessitating further data sampling. Inspired by OpenThoughts (guha2025openthoughts), we employ a difficulty-based weighted stratified sampling strategy on STEM-related data to achieve a balance between reasoning density and data diversity. Specifically, token length of the response is adopted as a natural and effective proxy for question difficulty. Our subsequent experiment corroborates the finding of OpenThoughts that length-based sampling outperforms annotation-based alternatives.
@@ 3.3 The curated Dataset Statistics
Figure 4: Comparison of source dataset distribution and curation statistics. Nemotron represents for nvidia-Llama-Nemotron-Post-Training-Dataset (bercovich2025llama), NatReason for naturalReasoning (yuan2025naturalreasoning), OSR-2 for OpenScienceReasoning-2.
@@ 4 Failure-Driven Post-training via Knowledge Enhancement
The widely used post-training pipeline (ouyang2022training) consists of two stages: SFT followed by second-stage RL. From our distribution-matching perspective (section˜2), the second stage aims to better approximate the target distribution, and can be implemented via either SFT or RL. Accordingly, in this section we introduce a failure-driven post-training pipeline that is both effective and efficient. Our paradigm is theoretically grounded to better fit the gold-standard distribution for reasoning ability, and is sample-efficient as it correctly leans towards the regions that contribute most to distribution matching.
@@ 4.1.2 The Failure-driven Second-stage Training
After the first-stage SFT and obtaining $\theta_{1}$, the model has already gained the basic information. To find the regions (i.e., query-response pairs) that the model underestimated (high density ratio between $P^{*}$ and $P_{0}$) or highly biased (high sample-wise loss $\ell_{\theta}(x,y)$), we evaluate the model on a distribution constructed from gold-standard evaluation tasks (such as AIME2025 (AIME2025), MMLU (wang2024mmlu) etc.) denoted as a distribution $Q$, and collect failure cases from this evaluation distribution. We define a failure function $w_{\theta_{1}}(x)\in\mathbb{R}_{+}$ to evaluate the value of the current sample to the model training, which in our implementation is a binary indicator of verified incorrectness:
@@
Given a retrieved document instance $d$, we synthesize a new training example $(x^{\prime},y^{\prime})$ via powerful LLMs such as DeepSeek R1 (guo2025deepseek). We view synthesis as another conditional kernel $G((x^{\prime},y^{\prime})\mid d)$ and apply an acceptance filter $a(x^{\prime},y^{\prime})\in\{0,1\}$ (e.g., answer-consistency verification), yielding $G_{\mathrm{acc}}(x^{\prime},y^{\prime}\mid d)\propto G(x^{\prime},y^{\prime}\mid d)\,a(x^{\prime},y^{\prime})$.
@@ 4.1.3 Extension to Reinforcement learning
Practically, we test our model using GRPO (shao2024deepseekmathpushinglimitsmathematical) and DAPO (yu2025dapoopensourcellmreinforcement), which we defer the introduction in section˜4.2.4.
@@ 4.2.2 Knowledge Retrieval
We first evaluate the SFT model on multiple established STEM benchmarks, including AIME2025 (AIME2025), MATH500, MMLU (wang2024mmlu) that works as the gold-standard distribution $Q$. We then collect questions answered incorrectly as “knowledge deficiency” samples. These erroneous instances are then encoded into dense vector representations using the Qwen3-8B-Embed [^6] https://huggingface.co/Qwen/Qwen3-Embedding-8B. Concurrently, every document in the knowledge base is embedded using the same model. Leveraging a vector retrieval algorithm, we retrieve the top-30 semantically most relevant documents for each incorrect question, ensuring high topical, conceptual, or methodological alignment between the query and retrieved content.
@@ 4.2.3 Data Synthesis
Utilizing the DeepSeek-R1 (guo2025deepseek) as the synthesis kernel $G((x^{\prime},y^{\prime})\mid d)$, we automatically generate two Query-Response pairs per retrieved document, each requiring deep reasoning based on the document’s core knowledge points and logical structure. To guarantee answer accuracy and consistency, a dual-verification mechanism is implemented to serve as the acceptance filter:
@@ 4.2.4 Reinforcement Learning Implementation
We test GRPO (shao2024deepseekmathpushinglimitsmathematical) and DAPO (yu2025dapo) for our framework, which defines the policy ratio $r_{i,t}$ and the renormalized advantage function $\hat{A}_{i}$ as,
@@
We use the open-source framework ROLL (wang2025reinforcementlearningoptimizationlargescale) to compute verified rewards. For math problems, we use math-verify to compare model outputs against ground-truth answers and assign a binary reward as advantage $A_{i}\in\{0,1\}$. For multiple-choice questions, we extract the option in $\texttt{\textbackslash boxed}\{$...\} and match it to the ground truth. Following yu2025dapoopensourcellmreinforcement, we apply dynamic sampling to filter problematic training instances by (i) dropping over-length responses and (ii) excluding prompts whose response-group mean scores fall outside a preset range. Hyperparameters are summarized in Table 12.
@@ 5.1.1 Benchmarks
Models are evaluated on representative reasoning benchmarks spanning mathematics and broader STEM disciplines. For mathematical reasoning, we assess performance on AIME2024 (AIME2024), AIME2025(AIME2025), HMMT_Feb_2025, BRUMO2025, BeyondAIME(beyondaime), and MATH-500. For remaining STEM domains, we adopt GPQA-Diamond (rein2024gpqa), R-Bench (rbench), MMLU-Pro (wang2024mmlu), and CMMLU (li2023cmmlu) as benchmarks. Particularly for MMLU-Pro and CMMLU, we select STEM-related subsets to specifically evaluate the model’s reasoning capabilities in STEM fields, as detailed in section˜A.4.
@@ 5.2.1 Sampling Strategy
Length-based sampling has been demonstrated to be an effective sampling method (guha2025openthoughts); nevertheless, our experiments demonstrate that its exclusive use degrades the model’s elementary reasoning and generalization capacities, as reported in table˜4. the ablation model trained solely with length-based sampling incurs significant performance drops on MATH500, MMLU-Pro-Math, and GPQA-Diamond, despite notable gains on AIME2024, AIME2025, and HMMT2025.
@@ C.3 Extension to RL: Theory
Let $\pi_{\theta}(y\mid x)$ denote the current policy and $\pi_{0}(y\mid x)$ be a reference policy (the stage-1 SFT model), which serves as a strong proposal distribution with broad support. Consider the following KL-regularized RL objective (DBLP:conf/icml/0011C00LZ23):

### v2

@@ 1 Introduction
In the past years, large language models (LLMs) like OpenAI’s o1 series [jaech2024openai], QwQ [qwq32b], and DeepSeek-R1 [guo2025deepseek] have demonstrated strong performance on challenging reasoning tasks in the field of mathematics and broader STEM. The reasoning capabilities in these models typically emerge from post-training techniques such as supervised fine-tuning (SFT) and/or reinforcement learning (RL) based on strong foundation models. However, while many models are open-sourced, the underlying post-training pipeline and training data curation remain undisclosed, creating both challenges and opportunities for future work.
More recently, the open-source community has made substantial efforts to develop data-construction recipes and algorithmic strategies for cultivating advanced reasoning capabilities in small-scale models, leading to a series of notable works such as Klear-Reasoner [su2025klear], Ring-Lite [DBLP:journals/corr/abs-2506-14731], MiMo [coreteam2025mimounlockingreasoningpotential], OpenThoughts [guha2025openthoughts], Llama-Nemotron [bercovich2025llama], and AceReason-Nemotron liu2025acereason, among others [fu2025areal, deepscaler2025, DBLP:journals/corr/abs-2503-10460]. However, despite these empirical successes, the community still lacks a unified framework to guide data curation and exploitation through post-training algorithms. It is widely recognized in the LLM community that “data is the new oil”, and the algorithm can only succeed when it effectively captures the desired data distribution. This motivates a central question in training reasoning models:
@@
In this report, we address this question from both theoretical and engineering perspectives. We first provide a data-centric view of the widely adopted SFT-RL pipeline [ouyang2022training] by framing it as a distribution-matching problem. We hypothesize that the first-stage SFT builds a strong proposal distribution that one draws samples for following usage; and the second stage post-training, no matter SFT or RL, shifts the model towards a gold-standard target distribution behind desirable properties such as reasoning ability.
@@
Specifically, we curate reasoning data from publicly available datasets and further augment it with synthetic examples generated from documents parsed by Logics-Parsing [chen2025logicsparsingtechnicalreport]. Coupled with a fine-grained, difficulty-aware stratified sampling strategy, extensive experiments show that our curated Logics-STEM-SFT-Dataset already equips LLMs with strong foundational reasoning capabilities. Moreover, we adopt a failure-driven post-training paradigm to further align the model with the gold-standard reasoning distribution. Concretely, after the first-stage SFT, we perform targeted knowledge retrieval and synthesize data around the model’s failure regions to guide a second-stage SFT or RL. This yields two alternative pipelines: SFT–RL and SFT–SFT. We systematically test our method under both SFT-RL and SFT-SFT pipelines and show that our approach substantially improves the model’s reasoning ability.
@@ 2 SFT-RL pipeline as a distribution matching
After pre-training, SFT followed by RL has become a widely adopted recipe for improving LLMs’ reasoning ability [ouyang2022training]. SFT is typically used to familiarize the model with long chain-of-thought (CoT) reasoning traces, while RL further aligns the model with human preferences or sharpens the policy distribution to produce more satisfactory responses with fewer samples [yue2025does, wen2025reinforcementlearningverifiablerewards]. From this perspective, the overall post-training procedure can be viewed as optimizing an expected objective estimated via Monte Carlo sampling. Let the training data consist of pairs $(x,y)$, where $x$ is the prompt and $y$ is the target output (e.g., a long-CoT response ending with a boxed final answer). Denote by $y^{t}$ the $t$-th token in yy [^1] We use subscripts $y_{i}$ to denote the $i$-th training example and superscripts $y^{t}$ to denote the $t$-th token within an example.. We write the sample-level supervised loss as $\ell_{\theta}(x,y)$ (e.g., the negative log-likelihood, NLL), where $\theta$ are the model parameters. The training objective is then,
@@
We elaborate on the remark a bit. SFT fits the model to $P_{0}$ through a large-scale dataset ${\mathcal{D}}\sim P_{0}$, which provides good proposal distribution with broad coverage and grants the model with general reasoning abilities. However, at this stage, the mismatch between $P_{0}$ and $P^{*}$ would make the optimization biased, thus performing suboptimally on specific tasks. The second stage training is used to implicitly eliminate the bias. For instance, a vanilla policy gradient can be considered as fitting a distribution depending on the advantage function $A(x,y)$ [DBLP:conf/icml/0011C00LZ23],
@@ 3.1 Data Curation Pipeline
Previous studies [guha2025openthoughts, liu2025acereason, openr1] have demonstrated the effectiveness of open-source data. To fully leverage open-source data and ensure diversity, we aggregate questions from a wide range of highly regarded and frequently cited sources on Huggingface [^2] https://huggingface.co/, as listed in Appendix A.1. For the purpose of preventing potential quality risks, all early synthetic subsets and multi-modal subsets are excluded. Additional, we utilize DEGISNER [liu2025designer] to incorporate synthetic datasets from books and webpages (DLR-Book and DLR-Web in liu2025designer) to further scale dataset volume and enhance diversity. All PDF-format documents are parsed into plain text by Logics-Parsing [chen2025logicsparsingtechnicalreport] and subsequently processed through the DESIGNER synthesis pipeline.
@@
To balance computational efficiency and response accuracy, Qwen3-235B-A22B-Thinking-2507444https://huggingface.co/Qwen/Qwen3-235B-A22B-Thinking-2507 serves as the teacher model to distill reasoning responses for each question. Generation configurations are detailed in Appendix A.3. To suppress systematic bias during subsequent training, we discard responses that exhibit degenerative repetition. Concretely, any response with an n-gram duplication ratio exceeding a predefined threshold is excluded from the training dataset. Optional further answer verification is conducted using math-verify library [^5] https://github.com/huggingface/Math-Verify. In cases where responses do not match the standard verifiable answers, the teacher model is employed to regenerate the response once more. The strategy of filtering out incorrect responses is not adopted, as empirical results indicate that it leads to degraded performance, consistent with the conclusion of OpenThoughts [guha2025openthoughts].
@@ 3.2 Weighted Stratified Sampling
Despite the implementation of extensive filtering strategies, the volume of data generated by our pipeline remains a significant challenge for model training, thereby necessitating further data sampling. Inspired by OpenThoughts [guha2025openthoughts], we employ a difficulty-based weighted stratified sampling strategy on STEM-related data to achieve a balance between reasoning density and data diversity. Specifically, token length of the response is adopted as a natural and effective proxy for question difficulty. Our subsequent experiment corroborates the finding of OpenThoughts that length-based sampling outperforms annotation-based alternatives.
@@ 3.3 The curated Dataset Statistics
Figure 4: Comparison of source dataset distribution and curation statistics. Nemotron represents for nvidia-Llama-Nemotron-Post-Training-Dataset [bercovich2025llama], NatReason for naturalReasoning [yuan2025naturalreasoning], OSR-2 for OpenScienceReasoning-2.
@@ 4 Failure-Driven Post-training via Knowledge Enhancement
The widely used post-training pipeline [ouyang2022training] consists of two stages: SFT followed by second-stage RL. From our distribution-matching perspective (section˜2), the second stage aims to better approximate the target distribution, and can be implemented via either SFT or RL. Accordingly, in this section we introduce a failure-driven post-training pipeline that is both effective and efficient. Our paradigm is theoretically grounded to better fit the gold-standard distribution for reasoning ability, and is sample-efficient as it correctly leans towards the regions that contribute most to distribution matching.
@@ 4.1.2 The Failure-driven Second-stage Training
After the first-stage SFT and obtaining $\theta_{1}$, the model has already gained the basic information. To find the regions (i.e., query-response pairs) that the model underestimated (high density ratio between $P^{*}$ and $P_{0}$) or highly biased (high sample-wise loss $\ell_{\theta}(x,y)$), we evaluate the model on a distribution constructed from gold-standard evaluation tasks (such as AIME2025 [AIME2025], MMLU [wang2024mmlu] etc.) denoted as a distribution $Q$, and collect failure cases from this evaluation distribution. We define a failure function $w_{\theta_{1}}(x)\in\mathbb{R}_{+}$ to evaluate the value of the current sample to the model training, which in our implementation is a binary indicator of verified incorrectness:
@@
Given a retrieved document instance $d$, we synthesize a new training example $(x^{\prime},y^{\prime})$ via powerful LLMs such as DeepSeek R1 [guo2025deepseek]. We view synthesis as another conditional kernel $G((x^{\prime},y^{\prime})\mid d)$ and apply an acceptance filter $a(x^{\prime},y^{\prime})\in\{0,1\}$ (e.g., answer-consistency verification), yielding $G_{\mathrm{acc}}(x^{\prime},y^{\prime}\mid d)\propto G(x^{\prime},y^{\prime}\mid d)\,a(x^{\prime},y^{\prime})$.
@@ 4.1.3 Extension to Reinforcement learning
Practically, we test our model using GRPO [shao2024deepseekmathpushinglimitsmathematical] and DAPO [yu2025dapoopensourcellmreinforcement], which we defer the introduction in section˜4.2.4.
@@ 4.2.2 Knowledge Retrieval
We first evaluate the SFT model on multiple established STEM benchmarks, including AIME2025 [AIME2025], MATH500, MMLU [wang2024mmlu] that works as the gold-standard distribution $Q$. We then collect questions answered incorrectly as “knowledge deficiency” samples. These erroneous instances are then encoded into dense vector representations using the Qwen3-8B-Embed [^6] https://huggingface.co/Qwen/Qwen3-Embedding-8B. Concurrently, every document in the knowledge base is embedded using the same model. Leveraging a vector retrieval algorithm, we retrieve the top-30 semantically most relevant documents for each incorrect question, ensuring high topical, conceptual, or methodological alignment between the query and retrieved content.
@@ 4.2.3 Data Synthesis
Utilizing the DeepSeek-R1 [guo2025deepseek] as the synthesis kernel $G((x^{\prime},y^{\prime})\mid d)$, we automatically generate two Query-Response pairs per retrieved document, each requiring deep reasoning based on the document’s core knowledge points and logical structure. To guarantee answer accuracy and consistency, a dual-verification mechanism is implemented to serve as the acceptance filter:
@@ 4.2.4 Reinforcement Learning Implementation
We test GRPO [shao2024deepseekmathpushinglimitsmathematical] and DAPO [yu2025dapo] for our framework, which defines the policy ratio $r_{i,t}$ and the renormalized advantage function $\hat{A}_{i}$ as,
@@
We use the open-source framework ROLL [wang2025reinforcementlearningoptimizationlargescale] to compute verified rewards. For math problems, we use math-verify to compare model outputs against ground-truth answers and assign a binary reward as advantage $A_{i}\in\{0,1\}$. For multiple-choice questions, we extract the option in $\texttt{\textbackslash boxed}\{$...\} and match it to the ground truth. Following yu2025dapoopensourcellmreinforcement, we apply dynamic sampling to filter problematic training instances by (i) dropping over-length responses and (ii) excluding prompts whose response-group mean scores fall outside a preset range. Hyperparameters are summarized in Table 12.
@@ 5.1.1 Benchmarks
Models are evaluated on representative reasoning benchmarks spanning mathematics and broader STEM disciplines. For mathematical reasoning, we assess performance on AIME2024 [AIME2024], AIME2025[AIME2025], HMMT_Feb_2025, BRUMO2025, BeyondAIME[beyondaime], and MATH-500. For remaining STEM domains, we adopt GPQA-Diamond [rein2024gpqa], R-Bench [rbench], MMLU-Pro [wang2024mmlu], and CMMLU [li2023cmmlu] as benchmarks. Particularly for MMLU-Pro and CMMLU, we select STEM-related subsets to specifically evaluate the model’s reasoning capabilities in STEM fields, as detailed in section˜A.4.
@@ 5.2.1 Sampling Strategy
Length-based sampling has been demonstrated to be an effective sampling method [guha2025openthoughts]; nevertheless, our experiments demonstrate that its exclusive use degrades the model’s elementary reasoning and generalization capacities, as reported in table˜4. the ablation model trained solely with length-based sampling incurs significant performance drops on MATH500, MMLU-Pro-Math, and GPQA-Diamond, despite notable gains on AIME2024, AIME2025, and HMMT2025.
@@ C.3 Extension to RL: Theory
Let $\pi_{\theta}(y\mid x)$ denote the current policy and $\pi_{0}(y\mid x)$ be a reference policy (the stage-1 SFT model), which serves as a strong proposal distribution with broad support. Consider the following KL-regularized RL objective [DBLP:conf/icml/0011C00LZ23]: